Calculate the final temperature that results from mixing 245 grams of cobalt, specific heat 0.446 J/g0C, at 1420C with 106 grams of water at 24.80C.




1
Expert's answer
2021-06-08T06:10:03-0400

Heat needed to raise water to 100"^0" C:

"0.106\times4.18\times(100-24.8)=.331Kj"

heat needed to change water to steam"= 2264.76kJ/kg\times0.106= 240kj"

Heat needed to raise steam from 100 to Tf.

"0.106\times2kj/kg\times(Tf-100)=???"


Now, heat given off by cooling Cobalt:

"0.245\times0.446\times(tf-1460)"

sums of heats = 0

"0.245\times0.446\times(tf-1460)-240-0.106\times2kj/kg\times(Tf-100)-0.331=0"

Now solve for "T_f" and easily get your answer.